The Cambridge Study in Delinquent Development was undertaken in north London (U.K.) to investigate the links between criminalWhat type of table is this? What is the variable presented in this table? How many boys had exactly 2 convictions by the endAre there outliers? (yes or no) Does the sample of boys used in this study represent a random sample of British boys?
Answer #1

1. The table is a frequency distribution table.

2. The variable presented in the table is Number of convictions

3. The number of boys who have exactly 2 conviction = Frequency corresponding to the number of convictions 2

= 21

4. Number of boys who had no convictions = 265

Total number of boys = 395

Percentage of boys who had no convictions = \frac{265}{395}*100=67.08\approx 67

ANS: The percentage of boys who had no conviction = 67%

5. The Frequency distribution is skewed. ( positively)

(as seen from the graph)

6. The frequency distribution is Unimodal. ( Here the mode { highest frequency } occurs only at 0 )

7.  Q1 is the 1/4*(N+1)th value = 395+1/4=396/4=99th value. where N = Total frequency =395

thus, Q1 = 0

Q3 is the 3/4*(N+1)th value = 3*(395+1)/4=3*396/4=297th value.

thus, Q3 = 1

Inter Quartile range =IQR= Q3-Q1=1-0=1

we can say that the values lying outside the range of (Q1-1.5*IQR,Q3+1.5*IQR) are potential outliers.

since (0-1.5*1,1+1.5*1)=(-1.5,2.5) So, the rest are detected potential outliers.

Hence YES there are outliers.

8. No, the sample doesnot represent a random sample.